how do i get y by itself in the following equations?
x+2y=4
2x+y=0
Please explain step by step thankyou

Answer by friesr(113)

x+2y=4
---
2y = -x +4
---
y = -x/2 + 2
===
2x+y=0
---
y = -2x
===